able building company is deciding whether to submiit a bid to build a new shopping centre. In the past, ABC's main competitor, Big Building Company (BBC) has submitted bids 70% of the time. If BBC does not bid for a job, the probability that ABC will get the job is 50%. If BBC bids on a job, the probability that ABC will get the job reduces to 25%. What is the probability that ABC will get the job?


1
Expert's answer
2021-06-01T03:47:52-0400

Let "A" be the event ''ABC will get the job'', "B" be the event '' BBC bids on a job''.

Given "P(B)=0.7. P(\bar{B})=1-0.7=0.3,"

"P(A|B)=0.25. P(A|\bar{B})=1-0.7=0.5"


The Law of Total Probability


"P(A)=P(A|B)P(B)+P(A|\bar{B})P(\bar{B})"

"=0.25(0.7)+0.5(0.3)=0.325"



The probability that ABC will get the job is 0.325.